Circular No. 7065

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ey IN NGC 7080 Ron Arbour, South Wonston, Hants, U.K., reports his discovery, on an

unfiltered CCD image taken on Dec. 5.816 UT with a 0.3-m f/6.3 Schmidt-Cassegrain reflector in the course of his supernova patrol, of an apparent supernova (mag ∼ 17.0) located at α = 21^{h}30^{m}00^{s}.56, δ = +26^{o}43′24′′.2 (equinox 2000.0), which is ∼ 18′′ west and 20′′ north of the center of NGC 7080. A prediscovery image of the new star appears on a frame taken by Arbour on Dec. 4.651, but it is not visible on frames taken Nov. 1 or earlier (limiting mag ∼ 17.5). The new object is not present in Vicker’s atlas or the second Digitized Sky Survey. M. Schwartz, Cottage Grove, OR, obtained an image of SN 1998ey at mag ∼ 16.5 on Dec. 8.108, using the Tenagra I telescope in Sonoita, AZ.

V4334 SAGITTARII G. Jacoby, Kitt Peak National Observatory (KPNO); and O. De Marco,

ETH Zentrum, Zurich, report: “The unusual variable star V4334 Sgr (cf. IAUC 6322) has continued to fade and become redder, further strengthening the argument for development of a dust shell or shells. Photometric measures from the KPNO 0.9-m telescope on 1998 June 4.27 UT are: V = 12.40, U − B = +1.07, B − V = +1.89, V − R = +1.38, V − I = +2.73. Observations by D. Willmarth at the WIYN 3.5-m telescope on Oct. 8.10 show a major drop in the optical, with V = 18.09, B − V = +2.89, V − R = +2.31, and V − I = +4.32. Additional WIYN data obtained by A. Saha and D. Harmer on Oct. 19.08 indicate a partial recovery to V = 16.91, B − V = +3.05, V − R = +2.11, and V − I = +3.77. The October data confirm reports on IAUC 7049 that V4334 Sgr had V > 15 during this period. Recoveries to V ∼ 15 in November (IAUC 7049) suggest a subsequent thinning of the October dust shell. Spectroscopic data taken at WIYN by D. Sawyer with the Densepak fiber-optic array on 1998 May 28.37 of the 420-nm CN band indicate a larger bandhead index as defined by Jacoby et al. (1998, A.J. 116, 1367) of 3.34 compared to an index of 2.32 during 1997 May and July. The increased band index is consistent with the cooler/redder atmosphere and/or the presence of a detached shell.”

1998 December 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green